OpenStudy (anonymous):

Computer sale ! Original Price:$598.00 25%off original price 8%tax applied after discount The owner offers a payment plan where the total cost of the computer is paid on 6months equal payments •determine the amount of each month payment

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ 25}}\%\ of\ 598 \implies \cfrac{{\color{brown}{ 25}}}{100}\cdot 598\implies {\color{brown}{ 25}}\div 100\cdot 598\implies \square \\ \quad \\ \textit{discounted price}\to 598-\square \\ \quad \\ {\color{brown}{ 8}}\%\ of\ (598-\square ) \implies \cfrac{{\color{brown}{ 8}}}{100}\cdot (598-\square )\implies {\color{brown}{ 8}}\div 100\cdot (598-\square ) \\ \quad \\ \textit{discounted price plus tax}\to (598-\square )\quad +\quad [{\color{brown}{ 8}}\div 100\cdot (598-\square )]\) to make it 6 even payments, divide it by 6